a point on he line x=3 from which tangents drawn to the circle x^2+y^2=8 are at right angles is ?

Grade:11

1 Answers

tarun
7 years ago
We know that pair of tangents from any pt fpr a circle are given by T^2=SS1.So when we write this equation for this case we put x1=3 as the pt is on x=3 and y coordinate as y1. We have a result that if a pair of straight lines are perpendicular a+b=0.applying this prop. We get y1=sq. root of 7.**don`t solve T^2=SS1 completely just obtain coeffs. Of x^2 and y^2.